1200 C of charge flow through a section of wire in 3 mins. How much current flows through it:

6.7 A

8.0 A

400 A

0.1 A

The "60W" label on a light bulb means

the bulb was filled with 60W at the factory

the bulb is expected to burn out after it uses up 60W

no matter how you use the bulb, it will supply 60 W of power

none of the other choices

Each of two small spheres is charged positively. The combined charge of the two spheres is 25 μC. Each sphere is repelled from the other by a force of magnitude 2 N when the two spheres are 20 cm apart. What is the charge on one of the spheres? Coulomb’s constant, k = 8.99 x 109 Nm2/C2.

29.64 μC

14.04 μC

24.64 μC

4.65 μC

Of the waves depicted below which one travels at the fastest in a vacuum?

A. Radio waves

B. Microwaves

C. Visible light

C

B

A

None

When a 1-Ω resistor, a 2-Ω resistor, and a 3-Ω resistor, are connected in series to a common DC power source, the least current will pass through

3-Ω resistor

None of them

2-Ω resistor

1-Ω resistor
Answer #1

Please comment if any doubts... will reply ASAP...

re O current = 9 - 1200 = 6674) + 3x60 2 Bulb expected burn after it uses 6ow ③ F = KQ, QL - 2 = k (Q.) (25-0,5*10*12 (0:2)2